Determine the amount of money in a savings account at the end of 3 years, given an initial deposit of $4,000 and a 4 percent annual interest rate when interest is compounded: Use Appendix A for an approximate answer, but calculate your final answer using the formula and financial calculator methods. (Do not round intermediate calculations. Round your final answers to 2 decimal places.)

a, annually

b. semiannually

c. quarterly
Answer #1
  • FV = Future Value
    PV = Present Value
    r = rate of interest
    n= no of period
    a) FV = PV (1 + r )n
    FV = 4000*(1+4%)^3
    FV = 4499.46
    b) FV = PV (1 + r )n
    FV = 4000*(1+4%/2)^6
    FV = 4504.65
    c) FV = PV (1 + r )n
    FV = 4000*(1+4%/4)^12
    FV = 4507.30